# Superfast Selection for Decision Tree Algorithms††thanks: Authors partially supported by NSF Grant IIS 1910131 and 1718945

### v1

@@ Abstract
We present a novel and systematic method, called Superfast Selection, for selecting the “optimal split" for decision tree and feature selection algorithms over tabular data. The method speeds up split selection on a single feature by lowering the time complexity, from O(MN) (using the standard selection methods) to O(M), where M represents the number of input examples and N the number of unique values. Additionally, the need for pre-encoding, such as one-hot or integer encoding, for feature value heterogeneity is eliminated. To demonstrate the efficiency of Superfast Selection, we empower the CART algorithm by integrating Superfast Selection into it, creating what we call Ultrafast Decision Tree (UDT). This enhancement enables UDT to complete the training process with a time complexity O(KMlogM) (K is the number of features). Additionally, the Training Only Once Tuning enables UDT to avoid the repetitive training process required to find the optimal hyper-parameter. Experiments show that the UDT can finish a single training on KDD99-10% dataset (494K examples with 41 features) within 1 second and tuning with 214.8 sets of hyper-parameters within 0.25 second on a laptop.
@@ 2 Superfast Selection Algorithm
where denotes the number of positive predictions and the number of negative predictions for the examples of with binary splitting, respectively; denotes the number of label classes, and denotes the number of examples. In Superfast Selection, this equation costs O() time to compute.
@@ 3 Ultrafast Decision Tree
Complexity: In the worst-case scenario for constructing a binary decision tree, the outcome is a balanced binary tree. The sorting of numerical values for K candidate features can be completed in O(KMlogM) time where M is the number of examples, and the sorting result can be reused throughout the entire tree-building process. The time complexity for building this tree is O(KMlogM), given that there are K candidate features and the cost for selecting a split on a single feature is O(M). Therefore, the Ultrafast Decision Tree algorithm powered by Superfast Selection can build decision trees within O(KMlogM) time in total.
@@ 5 Conclusions and Implications
In this paper, we present a highly efficient and systematic method of selecting the "optimal split" for decision tree algorithms, called Superfast Selection. Superfast Selection employs the prefix-sum technique to speed up the split selection on a single feature by lowering the time complexity to O(M) (M is the number of examples). Additionally, we presented our experiments with Ultrafast Decision Tree (UDT) that uses our Superfast Split method and which can complete training with a time complexity of O(KMlogM) (K is the number of features), and reported the results on 18 datasets. Furthermore, Superfast Selection empowers Ultrafast Decision Tree to handle hybrid features in a consolidated way eliminating the need of pre-encoding for feature-value heterogeneity. The result shows that even a lightweight laptop can finish training with UDT on datasets with millions of examples in seconds. Finally, Training Only Once Tuning helps UDT avoid repetitive training in the tuning process.

### v2

@@ Abstract
We present a novel and systematic method, called Superfast Selection, for selecting the “optimal split" for decision tree and feature selection algorithms over tabular data. The method speeds up split selection on a single feature by lowering the time complexity, from O(MN) (using the standard selection methods) to O(M), where M represents the number of input examples and N the number of unique values. Additionally, the need for pre-encoding, such as one-hot or integer encoding, for feature value heterogeneity is eliminated. To demonstrate the efficiency of Superfast Selection, we empower the CART algorithm by integrating Superfast Selection into it, creating what we call Ultrafast Decision Tree (UDT). This enhancement enables UDT to complete the training process with a time complexity O(KM2) (K is the number of features). Additionally, the Training Only Once Tuning enables UDT to avoid the repetitive training process required to find the optimal hyper-parameter. Experiments show that the UDT can finish a single training on KDD99-10% dataset (494K examples with 41 features) within 1 second and tuning with 214.8 sets of hyper-parameters within 0.25 second on a laptop.
@@ 2 Superfast Selection Algorithm
where denotes the number of positive predictions and the number of negative predictions for the examples of with binary splitting, respectively; C denotes the number of label classes, and M denotes the number of examples. In Superfast Selection, this equation costs O(C) time to compute.
@@ 3 Ultrafast Decision Tree
Complexity: Assuming the average case results in a balanced tree with a depth of O(logM). The sorting of numerical values for K candidate features can be completed in O(KMlogM) time where M is the number of examples, and the sorting result can be reused throughout the entire tree-building process. Therefore, the average bound of time complexity for building a UDT is (KMlogM), given that there are K candidate features and the cost for selecting a split on a single feature is O(M). In the worst-case scenario for constructing a binary decision tree, the outcome is an unbalanced tree with O(M) depth. The Ultrafast Decision Tree algorithm powered by Superfast Selection can build decision trees within O(KM2) time in total.
@@ 5 Conclusions and Implications
In this paper, we present a highly efficient and systematic method of selecting the "optimal split" for decision tree algorithms, called Superfast Selection. Superfast Selection employs the prefix-sum technique to speed up the split selection on a single feature by lowering the time complexity to O(M) (M is the number of examples). Additionally, we presented our experiments with Ultrafast Decision Tree (UDT) that uses our Superfast Split method and which can complete training with a time complexity of O(KM2) (K is the number of features), and reported the results on 18 datasets. Furthermore, Superfast Selection empowers Ultrafast Decision Tree to handle hybrid features in a consolidated way eliminating the need of pre-encoding for feature-value heterogeneity. The result shows that even a lightweight laptop can finish training with UDT on datasets with millions of examples in seconds. Finally, Training Only Once Tuning helps UDT avoid repetitive training in the tuning process.